OpenStudy (anonymous):

What is the diameter of the following circle: (x + 4)2 + (y – 9)2 = 18

OpenStudy (user917):

from the last equation (general equation of the circle) we can conclude that r^2=18 that mean r=sqrt(18) and then the diameter is twice r that means : diameter=2sqrt(18)=6sqrt(2).

OpenStudy (anonymous):

2 × √18 = 6√2 so the diameter would be = 8.485 ?

OpenStudy (anonymous):

Am I correct?